### v1

# Space-time FEM-BEM couplings for parabolic transmission problems

Abstract.
We develop couplings of a recent space-time first-order system least-squares (FOSLS) method for parabolic problems and space-time boundary element methods (BEM) for the heat equation to numerically solve a parabolic transmission problem on the full space and a finite time interval. In particular, we demonstrate coercivity of the couplings under certain restrictions and validate our theoretical findings by numerical experiments.
Key words and phrases:
parabolic transmission problem, space-time method, coupling, least-squares, boundary element method
2010 Mathematics Subject Classification:
35K20, 65M12, 65M60, 65N38
1. Introduction
1.1. Model problem
For $d\geq 2$, let $\Omega\subset\mathbb{R}^{d}$ be a bounded Lipschitz domain with connected boundary $\Gamma$ and $T>0$ a given final time with corresponding time interval $I:=(0,T)$. We abbreviate the space-time cylinder by $Q:=I\times\Omega$, and denote its lateral boundary by $\Sigma:=I\times\Gamma$. The corresponding outward normal vector on $\Sigma$ is denoted by ${\boldsymbol{n}}=(0,{\boldsymbol{n}}_{\boldsymbol{x}})\in\mathbb{R}^{d+1}$. Moreover, let $\Omega^{\rm ext}:=\mathbb{R}^{d}\setminus\overline{\Omega}$ and $Q^{\rm ext}:=I\times\Omega^{\rm ext}$ be the corresponding exterior domains. For ${\boldsymbol{A}}\in L^{\infty}(Q)^{d\times d}$ symmetric uniformly positive definite and given data $f:Q\to\mathbb{R}$, $u_{0}:\Omega\to\mathbb{R}$, $u_{D}:\Sigma\to\mathbb{R}$, and $\phi_{N}:\Sigma\to\mathbb{R}$, whose properties will be specified below, we consider the parabolic transmission problem to find $u:Q\rightarrow\mathbb{R}$ and $u^{\rm ext}:Q^{\rm ext}\rightarrow\mathbb{R}$ such that
[…]
Here, the convergence in (1.1g) is to be understood uniformly for $t\in I$ and ${\boldsymbol{x}}$ with $|{\boldsymbol{x}}|\to\infty$.
1.2. Motivation and existing works
Problem (1.1) models, for instance, heat conduction in a heterogeneous body $\Omega$ that radiates heat into the surrounding homogeneous medium $\Omega^{\rm ext}$. In this context, the unbounded domain $\Omega^{\rm ext}$ can of course be viewed as a simplification of a surrounding bounded $\widetilde{\Omega}$ that is substantially larger than $\Omega$. The work [42] provides a concrete interpretation of (1.1) in terms of photothermal spectroscopy. Other works that investigate the heat equation on unbounded domains include for instance [4, 44, 57]. Assuming that ${\boldsymbol{A}}$ is scalar and constant on $Q$, $f=0$ on $Q$, and $u_{0}=0$ on $\Omega$, [41] discretize (1.1) via a boundary element method (BEM) in the interior and the exterior, and the resulting semidiscrete system is solved by convolution quadrature [31, 33, 32] in time. In the PhD thesis [8], a coupling of the space-time finite element method (FEM) [50] based on some Hilbert-type transformation and space-time BEM [1, 39, 5] is proposed, and a numerical example for $d=1$ is presented. However, no numerical analysis of the scheme is given, and it is unclear whether the resulting discrete systems are inf-sup stable or even uniquely solvable. In addition, the computation of the mentioned Hilbert-type transformation is possible but non-trivial for higher dimensions [51, 54, 55]. A provably stable numerical scheme that approximately solves (1.1) in its general form has been missing in the literature up to now. It is the objective of this manuscript to introduce and analyze such a method.
A class of problems related to (1.1), so-called parabolic-elliptic transmission problems in which (1.1b) and (1.1g) for $u^{\rm ext}$ are replaced by the Laplace equation with appropriate radiation condition, have been intensely studied, e.g., in [34, 6, 5, 11, 12]. As a matter of fact, [34] models electromagnetic eddy-current problems as parabolic-hyperbolic transmission problems, where (1.1b) is actually replaced by a wave equation with small, potentially negligible, coefficient in front of the involved second time derivative of $u^{\rm ext}$.
For transmission problems with stationary elliptic PDEs in the interior and the exterior, FEM-BEM coupling is a well-understood numerical method, see, e.g., the overview article [2] and the references therein. We only mention the recent works [19, 18], which couple a discontinuous Petrov–Galerkin (DPG) method with standard as well as least-squares boundary element methods. Indeed, the couplings for the parabolic transmission problem (1.1) which we will introduce in the present manuscript were inspired by the couplings of those works.
1.3. Proposed couplings
We first reformulate the interior problem (1.1a)& (1.1c) as first-order system by substituting ${\boldsymbol{\sigma}}:=-{\boldsymbol{A}}\nabla_{\boldsymbol{x}}u$. For the heat equation, i.e., ${\boldsymbol{A}}={\bf Id}$, it has only recently been shown in [20] that the resulting first-order system least-squares (FOSLS) formulation, equipped with homogeneous Dirichlet boundary conditions for $u$, is well posed. This result was extended to general diffusion matrices ${\boldsymbol{A}}$ and inhomogeneous Dirichlet and/or Neumann boundary conditions in [24]. The results of [24] will be crucially exploited in the present work, but we also mention the recent generalizations [25, 17] to the instationary Stokes problem and the wave equation, respectively.
In a second step, we couple the FOSLS formulations [20, 24] of the interior problem (1.1a)&(1.1c) via the jump conditions (1.1e)&(1.1f) with boundary integral formulations [1, 39, 5] of the exterior problem (1.1b)&(1.1d)&(1.1g). While the exterior problem is posed on the unbounded $(d+1)$-dimensional set $Q^{\rm ext}$, its boundary integral reformulation is now posed on the bounded $d$-dimensional set $\Sigma$. We consider couplings of the FOSLS formulations to three different boundary integral formulations of the exterior problem: (i) the weakly-singular integral equation (with unknown Neumann data), (ii) the hypersingular integral equation (with unknown Dirichlet data), and (iii) the sum of the two preceding equations. Similar boundary integral formulations are also possible for the heat equation in the interior in case of homogeneous source term $f=0$ and homogeneous initial data $u_{0}=0$, which again yields a dimension reduction by one. Besides this dimension reduction, the space-time boundary integral equations of [1, 39, 5] have the advantage that the corresponding variational formulations are always coercive.
We prove coercivity of the resulting variational formulations of the three coupling methods, but we require the restriction ${\boldsymbol{A}}={\bf Id}$ for the first two of them. The reason for this is that we exploit a specific relation between the Dirichlet and Neumann data from the interior and the exterior problem in terms of boundary integral operators, which is only valid if we consider the same PDE in the interior and the exterior. In the stationary case, [45, 48] removed this restriction for the Johnson–Nédélec coupling, allowing for spatial diffusion matrices in the interior with minimal eigenvalue strictly larger than $1/4$. The latter can even be slightly improved, cf. [40]. A generalization to problem (1.1) and the FEM-BEM couplings considered here is not straight-forward. However, no such restriction is required for the third coupling involving all four boundary integral operators of the weakly-singular and the hypersingular equation. For all three coupling methods, the proof of coercivity requires the FOSLS part to be weighted with a sufficiently large positive number $\alpha>0$ whose size is in principle unknown. That being said, in our numerical experiments (for the first coupling), we always choose $\alpha=1$ without encountering any issues. We also present numerical experiments showing the effect of $\alpha$ on the coercivity of the resulting discrete systems. In these experiments on the unit square, only a small weighting of the order of $0.1$ or less results in a Galerkin matrix which is not positive definite.
1.4. Space-time discretization
The coupling methods presented in this work are coercive and fit thus into the setting of the Lax–Milgram lemma. Hence, in contrast to time-stepping methods as [41], they provide quasi-optimal approximations of the solution $(u,{\boldsymbol{\sigma}})=(u,-{\boldsymbol{A}}\nabla_{\boldsymbol{x}}u)$ for arbitrary discrete trial spaces. (The solution in the exterior $u^{\rm ext}$ can be computed in a post-processing step via a representation formula.) Regarding discretizations of the FOSLS methods [20, 24] for the interior problem with homogeneous Dirichlet boundary conditions, standard Lagrange elements on simplicial meshes of the full space-time cylinder $Q$ can lead to disappointingly low convergence rates, cf. [20]. However, [27] proposes discrete trial spaces on prismatic meshes with enhanced approximation properties, requiring less regularity of the unknown exact solution $(u,{\boldsymbol{\sigma}})$. This trial space will also be used in the present work.
The fact that arbitrary discrete trial spaces are admissible for our couplings allows for a further important advantage of space-time methods, namely the possibility of local mesh refinement simultaneously in both space and time. We intend to investigate suitable a posteriori error estimators to adaptively steer space-time local mesh refinement in the future. In this regard, we remark that the least-squares functional of the space-time FOSLS formulation [20, 24] automatically provides a reliable and efficient estimator, and that [28] recently introduced an estimator for the weakly-singular boundary integral equation. For a posteriori error analysis of stationary FEM-BEM couplings, we refer to [2] and the references therein.
While a time-stepping method as [41] only needs to solve relatively small linear systems in each time step, a potential downside of simultaneous space-time methods [1, 39, 5, 46, 49, 9, 50, 20, 24, 25, 17] is that one has to solve one huge linear system at once. In addition, boundary element discretizations lead to dense matrices due to the nonlocality of the boundary integral operators. On the other hand, simultaneous space-time methods in principle have the potential for massive parallelization, see, e.g., [23, 38, 52] for FEM and [10, 56] for BEM. Preconditioners for the employed space-time FOSLS formulation [20, 24] that can be applied in parallel are still missing in the literature and shall be addressed in the future. Having efficient preconditioners for FEM and BEM at hand, block-diagional preconditioners give efficient preconditioners for FEM-BEM couplings as has been analyzed in, e.g., [37, 15] for elliptic transmission problems. Moreover, the dense BEM matrices can be compressed via the fast multipole method and $\mathcal{H}$-matrices, cf. [35, 36, 29, 53].
Finally, coercivity also allows to easily treat optimal control and/or parameter-dependent problems for (1.1) as in the works [26, 21, 22] for the space-time FOSLS formulation of parabolic PDEs [20, 24]. Note that applications to optimal control particularly make the potential advantage of time-stepping methods requiring to store the solution only at each time step obsolete, as the solution on the full space-time cylinder is needed simultaneously for these problems anyway.
1.5. Outline
In Section 2.1, we briefly recall anisotropic Sobolev spaces on the (exterior) space-time cylinder and its lateral boundary. We further recall the definition and mapping properties of the corresponding boundary integral operators as well as the corresponding representation formula for solutions of the interior heat equation in Section 2.2. Section 2.3 then states the precise weak formulation of the considered parabolic transmission problem (1.1). In particular, a representation formula for the exterior solution $u^{\rm ext}$ is derived in terms of its Cauchy data, exploiting the radiation condition (1.1g). Section 3.1 briefly summarizes the employed results of [24] on space-time FOSLS formulations of second-order parabolic problems with inhomogeneous Dirichlet and Neumann boundary conditions. Section 3.2 and 3.3 constitute the core of the paper, introducing three different couplings of FOSLS for the interior problem with boundary integral equations for the exterior problem and proving their coercivity (under certain restrictions). In Section 3.4, we verify that these couplings are indeed equivalent to the transmission problem (1.1). We particularly show that (1.1) actually has a unique solution. The paper is concluded in Section 4 with numerical experiments, demonstrating the performance of the first coupling for smooth but also the limitations for non-smooth solutions under uniform mesh refinement. We also briefly comment on the implementation and analyze positive definiteness of the Galerkin matrices in dependence of the stabilization parameter $\alpha$ for the FOSLS part.
1.6. General notation
Throughout and without any ambiguity, ||$\cdot$|| ⋅ | denotes the absolute value of scalars, the Euclidean norm of vectors in $\mathbb{R}^{n}$, or the measure of a set in $\mathbb{R}^{n}$, e.g., the length of an interval or the area of a surface in $\mathbb{R}^{3}$. We write $A\lesssim B$ to abbreviate $A\leq CB$ with some generic constant $C>0$ which is clear from the context. Moreover, $A\eqsim B$ abbreviates $A\lesssim B\lesssim A$.
2. Preliminaries
2.1. Sobolev spaces
We use standard notation for Lebesgue and Sobolev spaces on bounded Lipschitz domains $\Omega$ and the corresponding boundary $\Gamma$, i.e., $L^{2}(\Omega)$, $H^{1}(\Omega)$, $H^{-1}(\Omega)=H^{1}_{0}(\Omega)^{\prime}$ and $H^{1/2}(\Gamma)$, $H^{-1/2}(\Gamma)=H^{1/2}(\Gamma)^{\prime}$. Similarly, we write $L^{2}(I;X)$ for the Bochner space of all measurable and square-integrable functions mapping from the time interval $I$ to some Banach space $X$. Moreover, we recall the definition of Bochner–Sobolev spaces $H^{\nu}(I;X):=\big{\{}v\in L^{2}(I;X)\,:\,\|v\|_{H^{\nu}(I;X)}<\infty\big{\}}$ for $\nu\in(0,1]$, associated with the norm
$\|v\|_{H^{\nu}(I;X)}^{2}:=\|v\|_{L^{2}(I;X)}^{2}+|v|_{H^{\nu}(I;X)}^{2},\quad|v|_{H^{\nu}(I;X)}^{2}:=\begin{cases}\int_{I}\int_{I}\frac{\|v(t)-v(s)\|_{X}^{2}}{|t-s|^{1+2\nu}}\,{\rm d}s\,{\rm d}t&\text{ if }\nu\in(0,1),\\ \|\partial_{t}v\|_{L^{2}(I;X)}^{2}&\text{ if }\nu=1,\end{cases}$
where $\partial_{t}$ denotes the (weak) time derivative. Finally, we define the anisotropic Sobolev spaces
$H^{1,1/2}(Q)$ $:=L^{2}(I;H^{1}(\Omega))\cap H^{1/2}(I;L^{2}(\Omega)),$
$H^{1/2,1/4}(\Sigma)$ $:=L^{2}(I;H^{1/2}(\Gamma))\cap H^{1/4}(I;L^{2}(\Gamma)).$
We denote by $H^{-1/2,-1/4}(\Sigma)$ the dual space of $H^{1/2,1/4}(\Sigma)$ with duality pairing $\langle\cdot\,,\,\cdot\rangle_{\Sigma}$ and interpret $L^{2}(\Sigma)$ as subspace of $H^{-1/2,-1/4}(\Sigma)$ via
$\langle u\,,\,\phi\rangle_{\Sigma}:=\langle\phi\,,\,u\rangle_{\Sigma}:=\int_{\Sigma}u(t,{\boldsymbol{x}})\phi(t,{\boldsymbol{x}})\,{\rm d}{\boldsymbol{x}}\,{\rm d}t\quad\text{for all }u\in H^{1/2,1/4}(\Sigma)\text{ and }\phi\in L^{2}(\Sigma).$
We recall the continuous embedding
$L^{2}(I;H^{1}(\Omega))\cap H^{1}(I;H^{-1}(\Omega))\hookrightarrow H^{1,1/2}(Q)\cap C^{0}(\overline{I};L^{2}(\Omega))$ (2.1)
from, e.g., [7, pages 480 & 494]. Next, we recall that the restriction $(\cdot)|_{\Sigma}$ is a bounded linear mapping from $H^{1,1/2}(Q)$ to $H^{1/2,1/4}(\Sigma)$; see, e.g., [30, Chapter 4, Theorem 2.1]. Finally, [5, Proposition 2.18] states that the normal derivative $\nabla_{\boldsymbol{x}}(\cdot)\cdot{\boldsymbol{n}}_{\boldsymbol{x}}$ is a bounded linear mapping from the space $\big{\{}u\in H^{1,1/2}(Q)\,:\,\partial_{t}u-\Delta_{\boldsymbol{x}}u\in L^{2}(Q)\big{\}}$ equipped with the canonical graph norm to $H^{-1/2,-1/4}(\Sigma)$.
For $R>0$, let $B_{R}(\boldsymbol{0}):=\big{\{}{\boldsymbol{x}}\in\mathbb{R}^{d}\,:\,|{\boldsymbol{x}}|<R\big{\}}$ with boundary $\Gamma_{R}:=\partial B_{R}(\boldsymbol{0})$. For $R>0$ such that $\overline{\Omega}\subset B_{R}(\boldsymbol{0})$, we abbreviate $\Omega^{\rm ext}_{R}:=\Omega^{\rm ext}\cap B_{R}(\boldsymbol{0})$, $Q_{R}^{\rm ext}:=I\times\Omega_{R}^{\rm ext}$, and $\Sigma_{R}:=I\times\Gamma_{R}$ with corresponding outward normal vector ${\boldsymbol{n}}_{R}=(0,{\boldsymbol{n}}_{R,{\boldsymbol{x}}})$. We define the Sobolev spaces on $Q_{R}^{\rm ext}$ and $\Sigma_{R}$ as for $Q$ and $\Sigma$.
2.2. Boundary integral operators
Define the heat kernel
$G(t,{\boldsymbol{x}}):=\begin{cases}\frac{1}{(4\pi t)^{d/2}}\,e^{-\frac{|{\boldsymbol{x}}|^{2}}{4t}}\quad&\text{for }(t,{\boldsymbol{x}})\in(0,\infty)\times\mathbb{R}^{d},\\ 0\quad&\text{else}.\end{cases}$
Note that for $(t,{\boldsymbol{x}})\neq(0,\boldsymbol{0})$, the function $G$ is smooth and satisfies the homogeneous heat equation. For sufficiently smooth functions $\phi$ and $u$ on the lateral boundary $\Sigma$, we define the single-layer potential
$(\widetilde{\mathcal{V}}\phi)(t,{\boldsymbol{x}})$ $:=\int_{\Sigma}G(t-s,{\boldsymbol{x}}-{\boldsymbol{y}})\phi(s,{\boldsymbol{y}})\,{\rm d}{\boldsymbol{y}}\,{\rm d}s\quad\text{for all }(t,{\boldsymbol{x}})\in Q\cup Q^{\rm ext}$ (2.2)
and the double-layer potential
$(\widetilde{\mathcal{K}}u)(t,{\boldsymbol{x}})$ $:=\int_{\Sigma}\nabla_{\boldsymbol{y}}G(t-s,{\boldsymbol{x}}-{\boldsymbol{y}})\cdot{\boldsymbol{n}}_{\boldsymbol{y}}\,u(s,{\boldsymbol{y}})\,{\rm d}{\boldsymbol{y}}\,{\rm d}s\quad\text{for all }(t,{\boldsymbol{x}})\in Q\cup Q^{\rm ext}.$ (2.3)
These potentials can be extended to bounded linear mappings $\widetilde{\mathcal{V}}:H^{-1/2,-1/4}(\Sigma)\to H^{1,1/2}(Q)$ and $\widetilde{\mathcal{K}}:H^{1/2,1/4}(\Sigma)\to H^{1,1/2}(Q)$. For any $\phi\in H^{-1/2,-1/4}(\Sigma)$ and $u\in H^{1/2,1/4}(\Sigma)$, $\widetilde{\mathcal{V}}\phi$ and $\widetilde{\mathcal{K}}u$ are indeed potentials in the sense that they satisfy the heat equation. This also readily implies that $\widetilde{\mathcal{V}}\phi,\widetilde{\mathcal{K}}u\in L^{2}(I;H^{1}(\Omega)\cap H^{1}(I;H^{-1}(\Omega))$. Finally, the restriction of $\widetilde{\mathcal{V}}\phi$ and $\widetilde{\mathcal{K}}u$ to $\{0\}\times\Omega$ is zero, i.e., $(\widetilde{\mathcal{V}}\phi)(0,\cdot)=(\widetilde{\mathcal{K}}u)(0,\cdot)=0$. Conversely, if $u\in L^{2}(I;H^{1}(\Omega))\cap H^{1}(I;H^{-1}(\Omega))$ is a weak solution of the heat equation $\partial_{t}u-\Delta_{\boldsymbol{x}}u=0$ with $u(0,\cdot)=0$, then there holds the representation formula
$u=\widetilde{\mathcal{V}}(\nabla_{\boldsymbol{x}}u\cdot{\boldsymbol{n}}_{\boldsymbol{x}})-\widetilde{\mathcal{K}}(u|_{\Sigma})\quad\text{on }Q.$ (2.4)
We further define the single-layer operator $\mathcal{V}$ and the double-layer operator $\mathcal{K}$ as in (2.2)–(2.3) with $(t,{\boldsymbol{x}})\in I\times\Sigma$ instead of $(t,{\boldsymbol{x}})\in Q$. These operators on the lateral boundary $\Sigma$ can be extended to bounded linear mappings $\mathcal{V}:H^{-1/2,-1/4}(\Sigma)\to H^{1/2,1/4}(\Sigma)$ and $\mathcal{K}:H^{1/2,1/4}(\Sigma)\to H^{1/2,1/4}(\Sigma)$. In addition, $\mathcal{V}$ is coercive, i.e.,
$\langle\mathcal{V}\phi\,,\,\phi\rangle_{\Sigma}\gtrsim\|\phi\|_{H^{-1/2,-1/4}(\Sigma)}^{2}\quad\text{for all }\phi\in H^{-1/2,-1/4}(\Sigma).$ (2.5)
The restriction of the single-layer potential from the interior $Q$ and from the exterior $Q^{\rm ext}$ to $\Sigma$ coincides with the single-layer operator, i.e.,
$\big{(}(\widetilde{\mathcal{V}}\phi)|_{Q}\big{)}|_{\Sigma}=\big{(}(\widetilde{\mathcal{V}}\phi)|_{Q^{\rm ext}}\big{)}|_{\Sigma}=\mathcal{V}\phi.$ (2.6)
The restriction of the double-layer potential from the interior $Q$ and from the exterior $Q^{\rm ext}$ to $\Sigma$ is given by
$\big{(}(\widetilde{\mathcal{K}}u)|_{Q}\big{)}|_{\Sigma}=(\mathcal{K}-\tfrac{1}{2})u\quad\text{and}\quad\big{(}(\widetilde{\mathcal{K}}u)|_{Q^{\rm ext}}\big{)}|_{\Sigma}=(\mathcal{K}+\tfrac{1}{2})u.$ (2.7)
Finally, for sufficiently smooth functions $\phi$ and $u$ on the lateral boundary $\Sigma$, we define the boundary integral operators
$(\mathcal{N}\phi)(t,{\boldsymbol{x}})$ $:=\int_{\Sigma}\nabla_{\boldsymbol{x}}G(t-s,{\boldsymbol{x}}-{\boldsymbol{y}})\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\,\phi(s,{\boldsymbol{y}})\,{\rm d}{\boldsymbol{y}}\,{\rm d}s\quad\text{for all }(t,{\boldsymbol{x}})\in\Sigma,$ (2.8)
$({\mathcal{W}}u)(t,{\boldsymbol{x}})$ $:=-\int_{\Sigma}\nabla_{\boldsymbol{x}}(\nabla_{\boldsymbol{y}}G(t-s,{\boldsymbol{x}}-{\boldsymbol{y}})\cdot{\boldsymbol{n}}_{\boldsymbol{y}})\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\,u(s,{\boldsymbol{y}})\,{\rm d}{\boldsymbol{y}}\,{\rm d}s\quad\text{for all }(t,{\boldsymbol{x}})\in\Sigma.$ (2.9)
These operators can be extended to bounded linear mappings $\mathcal{W}:H^{1/2,1/4}(\Sigma)\to H^{-1/2,-1/4}(\Sigma)$ and $\mathcal{N}:H^{-1/2,-1/4}(\Sigma)\to H^{1/2,1/4}(\Sigma)$. In addition, $\mathcal{W}$ is coercive, i.e.,
$\langle\mathcal{W}u\,,\,u\rangle_{\Sigma}\gtrsim\|u\|_{H^{1/2,1/4}(\Sigma)}^{2}\quad\text{for all }u\in H^{1/2,1/4}(\Sigma).$ (2.10)
As a matter of fact, [5, Theorem 3.11] even states the following coercivity
$\begin{split}(\phi,u)\begin{pmatrix}\mathcal{V}&-\mathcal{K}\\ \mathcal{N}&\mathcal{W}\end{pmatrix}\begin{pmatrix}\phi\\ u\end{pmatrix}\gtrsim\|\phi\|_{H^{-1/2,-1/4}(\Sigma)}^{2}+\|u\|_{H^{1/2,1/4}(\Sigma)}^{2}\\ \text{for all }(\phi,u)\in H^{-1/2,-1/4}(\Sigma)\times H^{1/2,1/4}(\Sigma),\end{split}$ (2.11)
which immediately yields (2.5) and (2.10). The interior and the exterior normal derivative of the single-layer potential is given by
$\nabla_{\boldsymbol{x}}\big{(}(\widetilde{\mathcal{V}}\phi)|_{Q}\big{)}\cdot{\boldsymbol{n}}_{x}=(\mathcal{N}+\tfrac{1}{2})\phi\quad\text{and}\quad\nabla_{\boldsymbol{x}}\big{(}(\widetilde{\mathcal{V}}\phi)|_{Q^{\rm ext}}\big{)}\cdot{\boldsymbol{n}}_{x}=(\mathcal{N}-\tfrac{1}{2})\phi.$ (2.12)
The interior and exterior normal derivative of the double-layer potential coincide with $-\mathcal{W}$, i.e.,
$\nabla_{\boldsymbol{x}}\big{(}(\widetilde{\mathcal{K}}u)|_{Q}\big{)}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}=\nabla_{\boldsymbol{x}}\big{(}(\widetilde{\mathcal{K}}u)|_{Q^{\rm ext}}\big{)}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}=-\mathcal{W}u.$ (2.13)
For details on proofs and further results, we refer to the seminal works [1, 39, 5].
2.3. Weak formulation
Given right-hand sides $f\in L^{2}(Q)$, $u_{0}\in L^{2}(\Omega)$, $u_{D}\in H^{1/2,1/4}(\Sigma)$, and $\phi_{N}\in H^{-1/2,-1/4}(\Sigma)$, we look for weak solutions of (1.1), i.e., $u\in L^{2}(I;H^{1}(\Omega))\cap H^{1}(I;H^{-1}(\Omega))$ and $u^{\rm ext}|_{Q^{\rm ext}_{R}}\in L^{2}(I;H^{1}(\Omega^{\rm ext}_{R}))$ for all $R>0$ such that $\overline{\Omega}\subset B_{R}(\boldsymbol{0})$. Note that $\partial_{t}u^{\rm ext}-\Delta_{\boldsymbol{x}}u^{\rm ext}=0$ readily yields the additional regularity $u^{\rm ext}|_{Q^{\rm ext}_{R}}\in L^{2}(I;H^{1}(\Omega^{\rm ext}_{R}))\cap H^{1}(I;H^{-1}(\Omega^{\rm ext}_{R}))$; see [5, Lemma 2.15]. By the properties of the involved Sobolev spaces mentioned in Section 2.1, the left-hand sides in (1.1) are indeed all well defined.
Let $R>0$ be sufficiently large such that $\overline{\Omega}\subset B_{R}(\boldsymbol{0})$. With the associated single- and double-layer potential $\widetilde{\mathcal{V}}_{R}$ and $\widetilde{\mathcal{K}}_{R}$ on $\Sigma_{R}$, the representation formula (2.4) applied to $Q^{\rm ext}_{R}$ instead of $Q$ yields that any weak solution of the exterior problem (1.1b)&(1.1d) is given by
$u^{\rm ext}=\widetilde{\mathcal{K}}(u^{\rm ext}|_{\Sigma})-\widetilde{\mathcal{V}}(\nabla_{\boldsymbol{x}}u^{\rm ext}\cdot{\boldsymbol{n}}_{\boldsymbol{x}})+\widetilde{\mathcal{V}}_{R}(\nabla_{\boldsymbol{x}}u^{\rm ext}\cdot{\boldsymbol{n}}_{R,{\boldsymbol{x}}})-\widetilde{\mathcal{K}}_{R}(u^{\rm ext}|_{\Sigma_{R}})\quad\text{on }Q_{R}^{\rm ext}.$ (2.14)
In what follows, we show that the radiation condition (1.1g) implies that the last two terms in (2.14) vanish as $R\to\infty$. We start with the following auxiliary result, which provides a crude estimate for the gradient $\nabla_{\boldsymbol{x}}u^{\rm ext}$.
Lemma 2.1.
Let $u^{\rm ext}$ be a weak solution of the exterior problem (1.1b)&(1.1d) with radiation condition (1.1g). Then, for all $\epsilon>0$ there exists a constant $C(d,\epsilon,T)>0$ such that
$|\nabla_{\boldsymbol{x}}u^{\rm ext}(t,{\boldsymbol{x}})|\leq C(d,\epsilon,T)\,t^{-1/2-\epsilon}\quad\text{for }t\in I\text{ as }|{\boldsymbol{x}}|\to\infty.$ (2.15)
Proof.
By parabolic regularity, e.g., [16, Theorem 1.58], the homogeneous heat equation (1.1b) implies that $u^{\rm ext}$ is even a smooth strong solution of $\partial_{t}u^{\rm ext}-\Delta_{\boldsymbol{x}}u^{\rm ext}=0$ on $Q^{\rm ext}$. In particular, each component of $\nabla_{\boldsymbol{x}}u^{\rm ext}$ satisfies the homogeneous heat equation strongly on $Q^{\rm ext}$. We may thus apply the mean-value property for the heat equation, e.g., [14, Section 2.3, Theorem 3] for each component: Let $R>0$ be sufficiently large such that $\overline{\Omega}\subset B_{R}(\boldsymbol{0})$ and $|u^{\rm ext}(t,{\boldsymbol{x}})|\leq 1$ for all $t\in I$ and all ${\boldsymbol{x}}\in\Omega^{\rm ext}$ with $|{\boldsymbol{x}}|\geq R$. Now, we fix coordinates $t\in I$ and ${\boldsymbol{x}}\in\Omega^{\rm ext}$ with $|{\boldsymbol{x}}|\geq 2R$. For $r>0$, we define the corresponding heat ball
$\begin{split}E(t,{\boldsymbol{x}};r)&:=\big{\{}(s,{\boldsymbol{y}})\in\mathbb{R}^{d+1}\,:\,G(t-s,{\boldsymbol{x}}-{\boldsymbol{y}})\geq r^{-d}\big{\}}\\ &=\big{\{}(s,{\boldsymbol{y}})\in\mathbb{R}^{d+1}\,:\,t-\frac{r^{2}}{4\pi}\leq s\leq t,|{\boldsymbol{x}}-{\boldsymbol{y}}|^{2}\leq 2d(t-s)\ln\frac{r^{2}}{4\pi(t-s)}\big{\}}.\end{split}$ (2.16)
For sufficiently small $r>0$ with $E(t,{\boldsymbol{x}};r)\subset Q^{\rm ext}$, it holds the mean-value property
$\nabla_{\boldsymbol{x}}u^{\rm ext}(t,{\boldsymbol{x}})=\frac{1}{4r^{d}}\int_{E(t,{\boldsymbol{x}};r)}\nabla_{\boldsymbol{x}}u^{\rm ext}(s,{\boldsymbol{y}})\frac{|{\boldsymbol{x}}-{\boldsymbol{y}}|^{2}}{(t-s)^{2}}\,{\rm d}{\boldsymbol{y}}\,{\rm d}s.$ (2.17)
Note that for the choice $r=\sqrt{t}$, the resulting term $2d(t-s)\ln\frac{t}{4\pi(t-s)}$ of (2.16) is uniformly bounded by a generic constant for $t\in I$ and $\tfrac{3}{4\pi}t\leq s\leq t$. Hence, we may assume that $R^{2}$ is also larger than this upper bound to guarantee that the closed ball $B({\boldsymbol{x}};t,s):=\big{\{}{\boldsymbol{y}}\in\mathbb{R}^{d}\,:\,|{\boldsymbol{x}}-{\boldsymbol{y}}|^{2}\leq 2d(t-s)\ln\frac{t}{4\pi(t-s)}\big{\}}$ is contained in $\Omega^{\rm ext}\setminus B_{R}(\boldsymbol{0})$, and thus $E(t,{\boldsymbol{x}};\sqrt{t})\subset Q^{\rm ext}$. The mean-value property (2.17) and integration by parts hence yield that
$\nabla_{\boldsymbol{x}}u^{\rm ext}(t,{\boldsymbol{x}})=\frac{1}{4t^{d/2}}\int_{\frac{3}{4\pi}t}^{t}\int_{B({\boldsymbol{x}};t,s)}\nabla_{\boldsymbol{y}}u^{\rm ext}(s,{\boldsymbol{y}})\frac{|{\boldsymbol{x}}-{\boldsymbol{y}}|^{2}}{(t-s)^{2}}\,{\rm d}{\boldsymbol{y}}\,{\rm d}s$
$=\frac{1}{4t^{d/2}}\Big{(}\int_{\frac{3}{4\pi}t}^{t}\int_{B({\boldsymbol{x}};t,s)}u^{\rm ext}(s,{\boldsymbol{y}})\frac{2({\boldsymbol{x}}-{\boldsymbol{y}})}{(t-s)^{2}}\,{\rm d}{\boldsymbol{y}}\,{\rm d}s+\int_{\frac{3}{4\pi}t}^{t}\int_{\partial B({\boldsymbol{x}};t,s)}u^{\rm ext}(s,{\boldsymbol{y}}){\boldsymbol{n}}_{B({\boldsymbol{x}};t,s)}({\boldsymbol{y}})\frac{|{\boldsymbol{x}}-{\boldsymbol{y}}|^{2}}{(t-s)^{2}}\,{\rm d}{\boldsymbol{y}}\,{\rm d}s\Big{)}.$
Here, ${\boldsymbol{n}}_{B({\boldsymbol{x}};t,s)}({\boldsymbol{y}})$ abbreviates the outward normal vector corresponding to $B({\boldsymbol{x}};t,s)$. Within both integration domains, we can further exploit that $|u^{\rm ext}(s,{\boldsymbol{y}})|\leq 1$. Together with the use of polar coordinates, we see that
$|\nabla_{\boldsymbol{x}}u^{\rm ext}(t,{\boldsymbol{x}})|\leq\frac{1}{4t^{d/2}}\Big{(}\int_{\frac{3}{4\pi}t}^{t}\int_{B({\boldsymbol{x}};t,s)}\frac{2|{\boldsymbol{x}}-{\boldsymbol{y}}|}{(t-s)^{2}}\,{\rm d}{\boldsymbol{y}}\,{\rm d}s+\int_{\frac{3}{4\pi}t}^{t}\int_{\partial B({\boldsymbol{x}};t,s)}\frac{|{\boldsymbol{x}}-{\boldsymbol{y}}|^{2}}{(t-s)^{2}}\,{\rm d}{\boldsymbol{y}}\,{\rm d}s\Big{)}$
$\lesssim\frac{1}{t^{d/2}}\int_{\frac{3}{4\pi}t}^{t}(t-s)^{(d-3)/2}\Big{(}\ln\frac{t}{4\pi(t-s)}\Big{)}^{(d+1)/2}\,{\rm d}s=\frac{1}{t^{d/2}}\int_{0}^{\frac{1}{4\pi}t}s^{(d-3)/2}\Big{(}\ln\frac{t}{4\pi s}\Big{)}^{(d+1)/2}\,{\rm d}s.$
If $d=3$, the last integral is just $\tfrac{1}{2\pi}t$, which concludes the proof in this case even for $\epsilon=0$. If $d=2$, we write $\ln\frac{t}{4\pi s}=\ln t-\ln 4\pi-\ln s\leq\ln T-\ln s\leq 2\max\{|\ln T|,|\ln s|\}$ and get that
$|\nabla_{\boldsymbol{x}}u^{\rm ext}(t,{\boldsymbol{x}})|\lesssim\frac{1}{t}\int_{0}^{\frac{1}{4\pi}t}s^{-1/2}\Big{(}|\ln T|^{3/2}+|\ln s|^{3/2}\Big{)}\,{\rm d}s\lesssim t^{-1/2-\epsilon}.$
This concludes the proof. ∎
Proposition 2.2.
Let $u^{\rm ext}$ be a weak solution of the exterior problem (1.1b)&(1.1d) with radiation condition (1.1g). Then, it satisfies the representation formula
$u^{\rm ext}=\widetilde{\mathcal{K}}(u^{\rm ext}|_{\Sigma})-\widetilde{\mathcal{V}}(\nabla_{\boldsymbol{x}}u^{\rm ext}\cdot{\boldsymbol{n}}_{\boldsymbol{x}})\quad\text{on }Q^{\rm ext}.$ (2.18)
Proof.
It remains to show that the last two terms in (2.14) tend to zero. To this end, let $(t,{\boldsymbol{x}})\in Q_{R}^{\rm ext}$ be arbitrary but fixed. The simple estimate $|G(s,{\boldsymbol{y}})|\lesssim s^{1/2}|{\boldsymbol{y}}|^{-d-1}$ for all $s\geq 0$, ${\boldsymbol{y}}\in\mathbb{R}^{d}$ (see, e.g., the proof of [5, Theorem 3.11]), the preceding lemma, and $|\Gamma_{R}|\simeq R^{d-1}$ show for sufficiently large $R>0$ and $20<\epsilon<1/$ 2 that
$\big{|}\big{(}\widetilde{\mathcal{V}}_{R}(\nabla_{\boldsymbol{x}}u^{\rm ext}\cdot{\boldsymbol{n}}_{R,{\boldsymbol{x}}})\big{)}(t,{\boldsymbol{x}})\big{|}$ $=\Big{|}\int_{\Sigma_{R}}G(t-s,{\boldsymbol{x}}-{\boldsymbol{y}})(\nabla_{\boldsymbol{y}}u^{\rm ext}(s,{\boldsymbol{y}})\cdot{\boldsymbol{y}}\,|{\boldsymbol{y}}|^{-1})\,{\rm d}{\boldsymbol{y}}\,{\rm d}s\Big{|}$
$\lesssim\int_{0}^{t}\int_{\Gamma_{R}}(t-s)^{1/2}|{\boldsymbol{x}}-{\boldsymbol{y}}|^{-d-1}s^{-1/2-\epsilon}\,{\rm d}{\boldsymbol{y}}\,{\rm d}s$
$\lesssim R^{-2}\int_{0}^{t}(t-s)^{1/2}s^{-1/2-\epsilon}\,{\rm d}{\boldsymbol{y}}\,{\rm d}s\lesssim R^{-2}\to 0\quad\text{as }R\to\infty.$
Using the radiation condition (1.1g) instead of Lemma 2.1, we similarly see that
$\big{|}\big{(}\widetilde{\mathcal{K}}_{R}(u^{\rm ext}|_{\Sigma_{R}})\big{)}(t,{\boldsymbol{x}})\big{|}$ $=\Big{|}\int_{\Sigma_{R}}\nabla_{\boldsymbol{y}}G(t-s,{\boldsymbol{x}}-{\boldsymbol{y}})\cdot{\boldsymbol{y}}\,|{\boldsymbol{y}}|^{-1}\,u^{\rm ext}(s,{\boldsymbol{y}})\,{\rm d}{\boldsymbol{y}}\,{\rm d}s\Big{|}$
$\lesssim\int_{0}^{t}\int_{\Gamma_{R}}\frac{|({\boldsymbol{x}}-{\boldsymbol{y}})\cdot{\boldsymbol{y}}|}{(t-s)^{d/2+1}|{\boldsymbol{y}}|}\,e^{-\frac{|{\boldsymbol{x}}-{\boldsymbol{y}}|^{2}}{4(t-s)}}\,{\rm d}{\boldsymbol{y}}\,{\rm d}s$
$\lesssim\int_{0}^{t}\int_{\Gamma_{R}}(t-s)^{-1/2}|{\boldsymbol{x}}-{\boldsymbol{y}}|^{-d}\,\,{\rm d}{\boldsymbol{y}}\,{\rm d}s$
$\lesssim R^{-1}\int_{0}^{t}(t-s)^{-1/2}\,{\rm d}s\lesssim R^{-1}\to 0\quad\text{as }R\to\infty.$
This concludes the proof. ∎
By the preceding proposition, once the weak solution $u$ of (1.1) is known, owing to the jump conditions (1.1e) and (1.1f), the corresponding $u^{\rm ext}$ can be computed as
$u^{\rm ext}=\widetilde{\mathcal{K}}(u|_{\Sigma}-u_{D})-\widetilde{\mathcal{V}}(({\boldsymbol{A}}\nabla_{\boldsymbol{x}}u)\cdot{\boldsymbol{n}}_{\boldsymbol{x}}-\phi_{N}).$ (2.19)
Conversely, in the proof of [5, Theorem 3.11], it is shown that any function of the form
[…]
Remark 2.3.
To be more precise, the cited [5, Theorem 3.11] states for $u^{\rm ext}$ defined as in (2.20) that $u^{\rm ext}(\cdot,{\boldsymbol{x}})={\mathcal{O}}(|x|^{-d})$ and $\nabla_{\boldsymbol{x}}u^{\rm ext}(\cdot,{\boldsymbol{x}})\cdot{\boldsymbol{x}}={\mathcal{O}}(|x|^{-d})$ as $|x|\to\infty$. Using the same argument, one even sees that $u^{\rm ext}(\cdot,{\boldsymbol{x}})={\mathcal{O}}(|x|^{-\mu})$ and $\nabla_{\boldsymbol{x}}u^{\rm ext}(\cdot,{\boldsymbol{x}})={\mathcal{O}}(|x|^{-\mu})$ as $|x|\to\infty$ for all fixed $\mu\geq 0$ (with hidden constants depending only on $d$, $\mu$, $u^{\rm ext}$, and $T$). The assumption (1.1g) that $u^{\rm ext}$ converges to zero as $|{\boldsymbol{x}}|\to\infty$ thus automatically yields a much stronger decay of $u^{\rm ext}$ and its spatial gradient $\nabla_{\boldsymbol{x}}u^{\rm ext}$. On the other hand, to guarantee that the last two terms in (2.14) converge to zero, it is clear from our proof that one could even allow for a certain growth of $u^{\rm ext}$ (and $\nabla_{\boldsymbol{x}}u^{\rm ext}$). Here and in the following, the minimal required radiation condition would be to explicitly assume that those two terms vanish as $R\to\infty$.
3. Coupling of FOSLS and boundary integral equations
3.1. First-order system least squares formulation for interior problem
We define the spaces
$U$ $:=\boldsymbol{H}(\operatorname{\mathrm{div}};Q)\cap(L^{2}(I;H^{1}(\Omega))\times L^{2}(Q)^{d}),$
$U_{D}$ :=$:=\big{\{}(u,{\boldsymbol{\sigma}})\in U\,:\,u|_{\Sigma}=0\big{\}},$
$U_{N}$ :=$:=\big{\{}(u,{\boldsymbol{\sigma}})\in U\,:\,{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}=0\big{\}},$
$L$ $:=L^{2}(Q)\times L^{2}(Q)^{d}\times L^{2}(\Omega).$
Here, $\operatorname{\mathrm{div}}$ denotes the time-space divergence $\operatorname{\mathrm{div}}(u,{\boldsymbol{\sigma}})=\partial_{t}u+\operatorname{\mathrm{div}}_{\boldsymbol{x}}{\boldsymbol{\sigma}}$ and $\boldsymbol{H}(\operatorname{\mathrm{div}};Q)=\big{\{}(u,{\boldsymbol{\sigma}})\in L^{2}(Q)^{1+d}\,:\,\operatorname{\mathrm{div}}(u,{\boldsymbol{\sigma}})\in L^{2}(Q)\big{\}}$. The spaces $U,U_{D},U_{N}$ are equipped with the canonical graph norm
$\|(u,{\boldsymbol{\sigma}})\|_{U}^{2}:=\|u\|_{L^{2}(Q)}^{2}+\|\nabla_{\boldsymbol{x}}u\|_{L^{2}(Q)}^{2}+\|{\boldsymbol{\sigma}}\|_{L^{2}(Q)}^{2}+\|\operatorname{\mathrm{div}}(u,{\boldsymbol{\sigma}})\|_{L^{2}(Q)}^{2}.$
We further define the bounded linear operators $\mathcal{G}:U\to L$, $\mathcal{G}_{D}:U\to L\times H^{1/2,1/4}(\Sigma)$, and $\mathcal{\mathcal{G}}_{N}:U\to L\times H^{-1/2,-1/4}(\Sigma)$ by
$\mathcal{G}(u,{\boldsymbol{\sigma}})$ $:=\big{(}\operatorname{\mathrm{div}}{\boldsymbol{u}},{\boldsymbol{A}}\nabla_{\boldsymbol{x}}u+{\boldsymbol{\sigma}},u(0,\cdot)\big{)},$
$\mathcal{G}_{D}(u,{\boldsymbol{\sigma}})$ $:=\big{(}\mathcal{G}(u,{\boldsymbol{\sigma}}),u|_{\Sigma}\big{)},$
$\mathcal{G}_{N}(u,{\boldsymbol{\sigma}})$ $:=\big{(}\mathcal{G}(u,{\boldsymbol{\sigma}}),{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\big{)}.$
Note that ${\boldsymbol{u}}\cdot{\boldsymbol{n}}={\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}$ for all ${\boldsymbol{u}}=(u,{\boldsymbol{\sigma}})\in U$. By [24, Lemma 2.2],
$(u,{\boldsymbol{\sigma}})\mapsto u\text{ is a bounded linear mapping from $U$ to $L^{2}(I;H^{1}(\Omega))\cap H^{1}(I;H^{-1}(\Omega))$},$ (3.1)
and (2.1) implies that $\mathcal{G}$ is indeed well defined and bounded. According to [24, Theorem 2.8], the latter two operators, $\mathcal{G}_{D}$ and $\mathcal{\mathcal{G}}_{N}$, are even bounded linear isomorphisms. Hence, $\mathcal{G}|_{U_{D}}$ and $\mathcal{G}|_{U_{N}}$ are bounded linear isomorphisms as well.
We finally note that any weak solution $u$ of the interior problem (1.1a)&(1.1c) solves
$\mathcal{G}{\boldsymbol{u}}=(f,\boldsymbol{0},u_{0})\quad\text{with }{\boldsymbol{u}}=(u,{\boldsymbol{\sigma}})=(u,-{\boldsymbol{A}}\nabla_{\boldsymbol{x}}u)\in U,$ (3.2)
or equivalently, for arbitrary fixed $\alpha>0$,
$b_{\alpha}({\boldsymbol{u}},{\boldsymbol{v}}):=\alpha\langle\mathcal{G}{\boldsymbol{u}}\,,\,\mathcal{G}{\boldsymbol{v}}\rangle_{L}=\alpha\langle(f,\boldsymbol{0},u_{0})\,,\,\mathcal{G}{\boldsymbol{v}}\rangle_{L}\quad\text{for all }{\boldsymbol{v}}=(v,{\boldsymbol{\tau}})\in U.$ (3.3)
Here, $\langle\cdot\,,\,\cdot\rangle_{L}$ denotes the scalar product on $L$.
3.2. Coupling to exterior problem via boundary integral equations
We introduce couplings to three different boundary integral equations obtained from the representation formula (2.19) for $u^{\rm ext}$. While the first two couplings only involve two of the four boundary integral operators but are only provably coercive in case of ${\boldsymbol{A}}={\bf Id}$, the third coupling involves all four boundary integral operators but is coercive for any symmetric uniformly positive definite ${\boldsymbol{A}}\in L^{\infty}(Q)^{d\times d}$.
3.2.1. Dirichlet trace of representation formula
If $u^{\rm ext}$ is the corresponding weak solution of the transmission problem (1.1), its representation (2.19) restricted to $\Sigma$ (using the identities (2.6)–(2.7)), the jump conditions (1.1e), and the fact that $({\boldsymbol{A}}\nabla_{\boldsymbol{x}}u)\cdot{\boldsymbol{n}}_{\boldsymbol{x}}=-{\boldsymbol{u}}\cdot{\boldsymbol{n}}=-{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}$ from (3.2) yield that
$\mathcal{V}(-{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}})+(\tfrac{1}{2}-\mathcal{K})(u|_{\Sigma})=\mathcal{V}\phi_{N}+(\tfrac{1}{2}-\mathcal{K})u_{D}.$ (3.4)
Since $\mathcal{G}_{N}(U)=L\times H^{-1/2,-1/4}(\Sigma)$, this is equivalent to the following variational formulation:
$c_{1}({\boldsymbol{u}},{\boldsymbol{v}}):=\langle\mathcal{V}({\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}})+(\mathcal{K}-\tfrac{1}{2})(u|_{\Sigma})\,,\,{\boldsymbol{\tau}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\rangle_{\Sigma}=\langle-\mathcal{V}\phi_{N}+(\mathcal{K}-\tfrac{1}{2})u_{D}\,,\,{\boldsymbol{\tau}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\rangle_{\Sigma}$
[…]
Combining (3.3) and (3.4), we conclude that any solution of (1.1) with ${\boldsymbol{u}}=(u,-{\boldsymbol{A}}\nabla_{\boldsymbol{x}}u)$ satisfies
$\begin{split}b_{\alpha}({\boldsymbol{u}},{\boldsymbol{v}})+c_{1}({\boldsymbol{u}},{\boldsymbol{v}})=\alpha\langle(f,\boldsymbol{0},u_{0})\,,\,\mathcal{G}{\boldsymbol{v}}\rangle_{L}+\langle-\mathcal{V}\phi_{N}+(\mathcal{K}-\tfrac{1}{2})u_{D}\,,\,{\boldsymbol{\tau}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\rangle_{\Sigma}\\ \text{for all }{\boldsymbol{v}}=(v,{\boldsymbol{\tau}})\in U.\end{split}$ (3.5)
By considering ${\boldsymbol{v}}\in U_{N}$ and recalling that $\mathcal{G}(U_{N})=L$, one sees that (3.5) conversely implies (3.2) and thus also (3.4). Note that the left-hand side and the right-hand side in (3.5) are linear and continuous with respect to inputs ${\boldsymbol{u}},{\boldsymbol{v}}\in U$.
3.2.2. Neumann trace of representation formula
Alternatively, one can consider the normal derivative of (2.19) and use the identities (2.12)–(2.13). Then, the jump condition (1.1f) and $({\boldsymbol{A}}\nabla_{\boldsymbol{x}}u)\cdot{\boldsymbol{n}}_{\boldsymbol{x}}=-{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}$ from (3.2) yield the equation
$\mathcal{W}(u|_{\Sigma})+(\mathcal{N}+\tfrac{1}{2})(-{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}})=\mathcal{W}u_{D}+(\mathcal{N}+\tfrac{1}{2})\phi_{N},$ (3.6)
or equivalently, owing to $\mathcal{G}_{D}(U)=L\times H^{1/2,1/4}(\Sigma)$,
$c_{2}({\boldsymbol{u}},{\boldsymbol{v}}):=\langle\mathcal{W}(u|_{\Sigma})+(\mathcal{N}+\tfrac{1}{2})(-{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}})\,,\,v|_{\Sigma}\rangle_{\Sigma}=\langle\mathcal{W}u_{D}+(\mathcal{N}+\tfrac{1}{2})\phi_{N}\,,\,v|_{\Sigma}\rangle_{\Sigma}$
for all $\text{for all }{\boldsymbol{v}}=(v,{\boldsymbol{\tau}})\in U.$
Combining (3.3) and (3.4), we conclude that any solution of (1.1) with ${\boldsymbol{u}}=(u,-{\boldsymbol{A}}\nabla_{\boldsymbol{x}}u)$ satisfies
$\begin{split}b_{\alpha}({\boldsymbol{u}},{\boldsymbol{v}})+c_{2}({\boldsymbol{u}},{\boldsymbol{v}})=\alpha\langle(f,\boldsymbol{0},u_{0})\,,\,\mathcal{G}{\boldsymbol{v}}\rangle_{L}+\langle\mathcal{W}u_{D}+(\mathcal{N}+\tfrac{1}{2})\phi_{N}\,,\,v|_{\Sigma}\rangle_{\Sigma}\\ \text{for all }{\boldsymbol{v}}=(v,{\boldsymbol{\tau}})\in U.\end{split}$ (3.7)
By considering ${\boldsymbol{v}}\in U_{D}$ and recalling that $\mathcal{G}(U_{D})=L$, one sees that (3.7) conversely implies (3.2) and thus also (3.6). Note that the left-hand side and the right-hand side in (3.7) are linear and continuous with respect to inputs ${\boldsymbol{u}},{\boldsymbol{v}}\in U$.
3.2.3. Dirichlet and Neumann trace of representation formula
Finally, one can also consider the sum of the two bilinear forms $c_{1}$ and $c_{2}$, which then corresponds to the sum of the two variational formulations, i.e.,
$c_{3}({\boldsymbol{u}},{\boldsymbol{v}}):=\langle\mathcal{V}({\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}})+(\mathcal{K}-\tfrac{1}{2})(u|_{\Sigma})\,,\,{\boldsymbol{\tau}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\rangle_{\Sigma}+\langle\mathcal{W}(u|_{\Sigma})+(\mathcal{N}+\tfrac{1}{2})(-{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}})\,,\,v|_{\Sigma}\rangle_{\Sigma}$
$=\langle-\mathcal{V}\phi_{N}+(\mathcal{K}-\tfrac{1}{2})u_{D}\,,\,{\boldsymbol{\tau}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\rangle_{\Sigma}+\langle\mathcal{W}u_{D}+(\mathcal{N}+\tfrac{1}{2})\phi_{N}\,,\,v|_{\Sigma}\rangle_{\Sigma}\quad\text{for all }{\boldsymbol{v}}=(v,{\boldsymbol{\tau}})\in U.$
We overall conclude from (1.1) that
$\begin{split}b_{\alpha}({\boldsymbol{u}},{\boldsymbol{v}})+c_{3}({\boldsymbol{u}},{\boldsymbol{v}})&=\alpha\langle(f,\boldsymbol{0},u_{0})\,,\,\mathcal{G}{\boldsymbol{v}}\rangle_{L}+\langle-\mathcal{V}\phi_{N}+(\mathcal{K}-\tfrac{1}{2})u_{D}\,,\,{\boldsymbol{\tau}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\rangle_{\Sigma}\\ &\quad+\langle\mathcal{W}u_{D}+(\mathcal{N}+\tfrac{1}{2})\phi_{N}\,,\,v|_{\Sigma}\rangle_{\Sigma}\quad\text{for all }{\boldsymbol{v}}=(v,{\boldsymbol{\tau}})\in U.\end{split}$ (3.8)
Contrary to Sections 3.2.1 and 3.2.2, it is not evident that (3.8) conversely implies (3.2) and/or either of the equations (3.4) and (3.6). However, in Section 3.4, we will see that this is indeed the case. Note that the left-hand side and the right-hand side in (3.8) are linear and continuous with respect to inputs ${\boldsymbol{u}},{\boldsymbol{v}}\in U$.
3.3. Coercivity of combined bilinear forms
Our next goal is to show that the continuous bilinear forms $b_{\alpha}+c_{i}:U\times U\to\mathbb{R}$ are coercive for $i\in\{1,2,3\}$ and sufficiently large $\alpha>0$. We start with the following auxiliary result.
Lemma 3.1.
It holds that
$\langle u|_{\Sigma}\,,\,-{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\rangle_{\Sigma}=\|{\boldsymbol{A}}^{1/2}\nabla_{\boldsymbol{x}}u\|_{L^{2}(Q)}^{2}+\tfrac{1}{2}\|u(T,\cdot)\|_{L^{2}(\Omega)}^{2}\gtrsim\|{\boldsymbol{u}}\|_{U}^{2}\quad\text{for all }{\boldsymbol{u}}=(u,{\boldsymbol{\sigma}})\in{\rm ker}\,\mathcal{G}.$
Proof.
The kernel of $\mathcal{G}$ is given by the set
${\rm ker}\,\mathcal{G}=\big{\{}(u,{\boldsymbol{\sigma}})\in U\,:\,\partial_{t}u-\operatorname{\mathrm{div}}_{\boldsymbol{x}}({\boldsymbol{A}}\nabla_{\boldsymbol{x}}u)=0\wedge{\boldsymbol{\sigma}}=-{\boldsymbol{A}}\nabla_{\boldsymbol{x}}u\wedge u(0,\cdot)=0\big{\}}.$
Let ${\boldsymbol{u}}=(u,{\boldsymbol{\sigma}})\in{\rm ker}\,\mathcal{G}$. Then, boundedness (3.1) shows that
$\|u\|_{L^{2}(I;H^{1}(\Omega))\cap H^{1}(I;H^{-1}(\Omega))}^{2}\eqsim\|{\boldsymbol{u}}\|_{U}^{2}\eqsim\|u\|_{L^{2}(Q)}^{2}+\|\nabla_{\boldsymbol{x}}u\|_{L^{2}(Q)}^{2}.$ (3.9)
Note that $({\boldsymbol{A}}\nabla_{\boldsymbol{x}}u)\cdot{\boldsymbol{n}}_{\boldsymbol{x}}=-{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}$. Integration by parts similar to [5, Equations (2.35)&(2.52)] together with $\partial_{t}u-\operatorname{\mathrm{div}}_{\boldsymbol{x}}({\boldsymbol{A}}\nabla_{\boldsymbol{x}}u)=0$ and $u(0,\cdot)=0$ gives that
$\langle u|_{\Sigma}\,,\,-{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\rangle_{\Sigma}$ $=\int_{Q}({\boldsymbol{A}}\nabla_{\boldsymbol{x}}u)\cdot\nabla_{\boldsymbol{x}}u\,{\rm d}{\boldsymbol{x}}\,{\rm d}t-\int_{Q}\big{(}\partial_{t}u-\operatorname{\mathrm{div}}_{\boldsymbol{x}}({\boldsymbol{A}}\nabla_{\boldsymbol{x}}u)\big{)}\,u\,{\rm d}{\boldsymbol{x}}\,{\rm d}t+\int_{Q}\partial_{t}u\,u\,{\rm d}{\boldsymbol{x}}\,{\rm d}t$
$=\|{\boldsymbol{A}}^{1/2}\nabla_{\boldsymbol{x}}u\|_{L^{2}(Q)}^{2}+\tfrac{1}{2}\|u(T,\cdot)\|_{L^{2}(\Omega)}^{2}.$
Next, we want to apply a standard compactness argument, more precisely, the Peetre–Tartar lemma [13, Lemma A.20]. In the notation from there (with $S$ instead of $T$, being already the end time point), we set
$X$ :=$:=\big{\{}v\in L^{2}(I;H^{1}(\Omega))\cap H^{1}(I;H^{-1}(\Omega))\,:\,\partial_{t}v-\operatorname{\mathrm{div}}_{\boldsymbol{x}}({\boldsymbol{A}}\nabla_{\boldsymbol{x}}v)=0\big{\}},$
$Y$ $:=L^{2}(Q)\times L^{2}(\Omega),$
$Z$ $:=L^{2}(Q),$
and
$A:X\to Y,\quad v\mapsto\big{(}{\boldsymbol{A}}^{1/2}\nabla_{\boldsymbol{x}}v,\tfrac{1}{\sqrt{2}}v(T,\cdot)\big{)},$
$S:X\to Z,\quad v\mapsto v.$
Again by (3.1), it holds that
$\|v\|_{X}\lesssim\|v\|_{L^{2}(Q)}+\|{\boldsymbol{A}}^{1/2}\nabla_{\boldsymbol{x}}v\|_{L^{2}(Q)}\leq\|Av\|_{Y}+\|Sv\|_{Z}\quad\text{for all }v\in X.$
Moreover, the operator $A$ is bounded owing to the continuous embedding (2.1). It is also injective: Indeed, if $Av=(0,0)$, then $\nabla_{\boldsymbol{x}}v=0$ and thus $\partial_{t}v=\operatorname{\mathrm{div}}_{\boldsymbol{x}}{\boldsymbol{A}}\nabla_{\boldsymbol{x}}v=0$ so that $v$ is constant. As $v(T,\cdot)=0$, we conclude that $v=0$. Moreover, the operator $S$ is compact. This follows from the continuous inclusion $X\hookrightarrow L^{2}(I;H^{1}(\Omega))\cap H^{1}(I;H^{-1}(\Omega))$ and the Aubin–Lions–Simons lemma [3, Theorem II.5.16] stating that the inclusion $L^{2}(I;H^{1}(\Omega))\cap H^{1}(I;H^{-1}(\Omega))$ $\hookrightarrow L^{2}(Q)$ is compact. Overall, the Peetre–Tartar lemma yields that
$\|v\|_{X}\lesssim\|Av\|_{Y}\quad\text{for all }v\in X.$
Since $\|Au\|_{Y}^{2}=\|{\boldsymbol{A}}^{1/2}\nabla_{\boldsymbol{x}}u\|_{L^{2}(Q)}^{2}+\tfrac{1}{2}\|u(T,\cdot)\|_{L^{2}(\Omega)}^{2}$, we conclude the proof with (3.9). ∎
For the proof of coercivity of the bilinear forms $c_{1}$ and $c_{2}$, we restrict to the case ${\boldsymbol{A}}={\bf Id}$.
Lemma 3.2.
Suppose that ${\boldsymbol{A}}={\bf Id}$. Then, the continuous bilinear forms $c_{1}:U\times U\to\mathbb{R}$ and $c_{2}:U\times U\to\mathbb{R}$ are coercive on the kernel of $\mathcal{G}$, i.e.,
$c_{i}({\boldsymbol{u}},{\boldsymbol{u}})\gtrsim\|{\boldsymbol{u}}\|_{U}^{2}\quad\text{for }i\in\{1,2\}\text{ and all }{\boldsymbol{u}}\in{\rm ker}\,\mathcal{G}.$
Proof.
Let ${\boldsymbol{u}}=(u,{\boldsymbol{\sigma}})\in{\rm ker}\,\mathcal{G}$. Since $\partial_{t}u-\Delta_{\boldsymbol{x}}u=0$ with $u(0,\cdot)=0$, the representation formula (2.4) restricted to $\Sigma$ and the identities (2.6)–(2.7) yield that
$\mathcal{V}(\nabla_{\boldsymbol{x}}u\cdot{\boldsymbol{n}}_{\boldsymbol{x}})+(\tfrac{1}{2}-\mathcal{K})(u|_{\Sigma})=u|_{\Sigma}.$
Note that $\nabla_{\boldsymbol{x}}u\cdot{\boldsymbol{n}}_{\boldsymbol{x}}=-{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}$. This yields that $c_{1}({\boldsymbol{u}},{\boldsymbol{u}})=\langle u|_{\Sigma}\,,\,-{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\rangle_{\Sigma}$. Similarly, taking the normal derivative of the representation formula (2.4) and the identities (2.12)–(2.13) yield that
$\mathcal{W}(u|_{\Sigma})+(\mathcal{N}+\tfrac{1}{2})(\nabla_{x}u\cdot{\boldsymbol{n}}_{\boldsymbol{x}})=\nabla_{x}u\cdot{\boldsymbol{n}}_{\boldsymbol{x}},$
and thus, $c_{2}({\boldsymbol{u}},{\boldsymbol{u}})=\langle u|_{\Sigma}\,,\,-{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\rangle_{\Sigma}$. In both cases, Lemma 3.1 concludes the proof. ∎
Obviously, the bilinear form $c_{3}$, being the sum of the two bilinear forms $c_{1}$ and $c_{2}$ is also coercive on the kernel of $\mathcal{G}$, provided that ${\boldsymbol{A}}={\bf Id}$. Using a different proof, it turns out that the bilinear form $c_{3}$ is coercive on the kernel of $\mathcal{G}$ even without any restrictions on the diffusion matrix ${\boldsymbol{A}}$ in the interior space-time cylinder $Q$.
Lemma 3.3.
The continuous bilinear form $c_{3}:U\times U\to\mathbb{R}$ is coercive on the kernel of $\mathcal{G}$, i.e.,
$c_{3}({\boldsymbol{u}},{\boldsymbol{u}})\gtrsim\|{\boldsymbol{u}}\|_{U}^{2}\quad\text{for all }{\boldsymbol{u}}\in{\rm ker}\,\mathcal{G}.$
Proof.
The coercivity (2.11) yields the existence of some constant $C>0$ such that for any ${\boldsymbol{u}}=(u,{\boldsymbol{\sigma}})\in U$,
$c_{3}({\boldsymbol{u}},{\boldsymbol{u}})$ $=\langle\mathcal{V}({\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}})+(\mathcal{K}-\tfrac{1}{2})(u|_{\Sigma})\,,\,{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\rangle_{\Sigma}+\langle\mathcal{W}(u|_{\Sigma})+(\mathcal{N}+\tfrac{1}{2})(-{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}})\,,\,u|_{\Sigma}\rangle_{\Sigma}$
$=(-{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}},u|_{\Sigma})\begin{pmatrix}\mathcal{V}&-\mathcal{K}\\ \mathcal{N}&\mathcal{W}\end{pmatrix}\begin{pmatrix}-{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\\ u|_{\Sigma}\end{pmatrix}+\langle-{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\,,\,u|_{\Sigma}\rangle_{\Sigma}$
$\geq C\big{(}\|{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\|_{H^{-1/2,-1/4}(\Sigma)}^{2}+\|u|_{\Sigma}\|_{H^{1/2,1/4}(\Sigma)}^{2}\big{)}+\langle-{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\,,\,u|_{\Sigma}\rangle_{\Sigma}$
$\geq\langle-{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\,,\,u|_{\Sigma}\rangle_{\Sigma}.$
Lemma 3.1 concludes the proof. ∎
With Lemma 3.2 and Lemma 3.3, coercivity of $b_{\alpha}+c_{i}$, $i\in\{1,2,3\}$, follows from the following abstract theorem.
Theorem 3.4.
Let $U$ and $L$ be Hilbert spaces and $\mathcal{G}:U\to L$ a continuous linear mapping such that $\mathcal{G}|_{U_{D}}$ is even a continuous linear isomorphism onto $L$ for some closed subspace $U_{D}\subseteq U$. For $\alpha>0$, define the continuous bilinear form $b_{\alpha}:U\times U$ by $b_{\alpha}({\boldsymbol{u}},{\boldsymbol{v}}):=\alpha\langle\mathcal{G}{\boldsymbol{u}}\,,\,\mathcal{G}{\boldsymbol{v}}\rangle_{L}$. Moreover, let $c:U\times U$ be a second continuous bilinear form that is coercive on ${\rm ker}\,\mathcal{G}$. Then, there exists $\alpha_{0}>0$ such that for all $\alpha\geq\alpha_{0}$, $b_{\alpha}+c$ is coercive on $U$, i.e.,
$b_{\alpha}({\boldsymbol{u}},{\boldsymbol{u}})+c({\boldsymbol{u}},{\boldsymbol{u}})\gtrsim\|{\boldsymbol{u}}\|_{U}^{2}\quad\text{for all }{\boldsymbol{u}}\in U.$
Proof.
The operator $\mathcal{G}$ induces the decomposition $U=U_{D}+{\rm ker}\,\mathcal{G}$: Indeed, each ${\boldsymbol{u}}\in U$ can be uniquely written as ${\boldsymbol{u}}=\mathcal{G}|_{U_{D}}^{-1}\mathcal{G}{\boldsymbol{u}}+({\boldsymbol{u}}-\mathcal{G}|_{U_{D}}^{-1}\mathcal{G}{\boldsymbol{u}})$. We abbreviate the corresponding projection with range $U_{D}$ and kernel ${\rm ker}\,\mathcal{G}$ by $\mathcal{P}:=\mathcal{G}|_{U_{D}}^{-1}\mathcal{G}$. Then, the assertion follows from the assumptions and the elementary calculations
$\|{\boldsymbol{u}}\|_{U}^{2}$ $\lesssim\|\mathcal{P}{\boldsymbol{u}}\|_{U}^{2}+\|{\boldsymbol{u}}-\mathcal{P}{\boldsymbol{u}}\|_{U}^{2}$
$\lesssim\|\mathcal{G}{\boldsymbol{u}}\|_{L}^{2}+c({\boldsymbol{u}}-\mathcal{P}{\boldsymbol{u}},{\boldsymbol{u}}-\mathcal{P}{\boldsymbol{u}})$
$=\|\mathcal{G}{\boldsymbol{u}}\|_{L}^{2}+c({\boldsymbol{u}},{\boldsymbol{u}})-c({\boldsymbol{u}},\mathcal{P}{\boldsymbol{u}})-c(\mathcal{P}{\boldsymbol{u}},{\boldsymbol{u}})+c(\mathcal{P}{\boldsymbol{u}},\mathcal{P}{\boldsymbol{u}})$
$\lesssim\|\mathcal{G}{\boldsymbol{u}}\|_{L}^{2}+c({\boldsymbol{u}},{\boldsymbol{u}})+\|{\boldsymbol{u}}\|_{U}\|\mathcal{P}{\boldsymbol{u}}\|_{U}+\|\mathcal{P}{\boldsymbol{u}}\|_{U}\|\mathcal{P}{\boldsymbol{u}}\|_{U}$
$\lesssim\|\mathcal{G}{\boldsymbol{u}}\|_{L}^{2}+c({\boldsymbol{u}},{\boldsymbol{u}})+\|{\boldsymbol{u}}\|_{U}\|\mathcal{G}{\boldsymbol{u}}\|_{L}$
$\leq(1+\tfrac{\delta^{-1}}{2})\|\mathcal{G}{\boldsymbol{u}}\|_{L}^{2}+c({\boldsymbol{u}},{\boldsymbol{u}})+\tfrac{\delta}{2}\|{\boldsymbol{u}}\|_{U}^{2}.$
Choosing $\delta>0$ sufficiently small, this proves coercivity for suitably chosen $\alpha_{0}$. ∎
Remark 3.5.
Obviously, coercivity of the bilinear form $c$ on the kernel of $\mathcal{G}$ is also necessary for coercivity of the bilinear form $b_{\alpha}+c$.
3.4. Equivalence of couplings to transmission problem
It is straight-forward to see that the first two couplings (3.5) and (3.7) are equivalent to the transmission problem (1.1), even without any restrictions on the diffusion matrix ${\boldsymbol{A}}$; see Theorem 3.7 and Remark 3.8 below. This is not evident for the third coupling (3.8), which involves all four boundary integral operators. However, we have already seen that (1.1) implies each of the couplings (3.5), (3.7), and (3.8), and that the latter are uniquely solvable provided that $\alpha$ is chosen sufficiently large (see Theorem 3.4) and ${\boldsymbol{A}}={\bf Id}$ in case of (3.5) and (3.7). If we knew that (1.1) has indeed a unique weak solution, the desired equivalence would thus follow at once. While the problem at hand has been considered before, we were unable to find a rigorous proof of unique solvability. Therefore, we address this seemingly open question in the following.
We introduce yet another (though practically unfeasible) coupling, starting from the first-order system (3.2) for ${\boldsymbol{u}}=(u,{\boldsymbol{\sigma}})\in U$ in the interior and the boundary integral equation (3.4), i.e., the representation (2.19) of the exterior solution $u^{\rm ext}$ restricted to $\Sigma$. Note that (3.4) is equivalent to
$-{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}+\mathcal{V}^{-1}(\tfrac{1}{2}-\mathcal{K})(u|_{\Sigma})=\phi_{N}+\mathcal{V}^{-1}(\tfrac{1}{2}-\mathcal{K})u_{D},$ (3.10)
or, owing to $\mathcal{G}_{D}(U)=L\times H^{1/2,1/4}(\Sigma)$,
$c_{4}({\boldsymbol{u}},{\boldsymbol{v}}):=\langle-{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}+\mathcal{V}^{-1}(\tfrac{1}{2}-\mathcal{K})(u|_{\Sigma})\,,\,v|_{\Sigma}\rangle_{\Sigma}=\langle\phi_{N}+\mathcal{V}^{-1}(\tfrac{1}{2}-\mathcal{K})u_{D}\,,\,v|_{\Sigma}\rangle_{\Sigma}$
for all $\text{for all }{\boldsymbol{v}}=(v,{\boldsymbol{\tau}})\in U.$
Combining (3.3) and (3.4), we conclude that any solution of (1.1) with ${\boldsymbol{u}}=(u,-{\boldsymbol{A}}\nabla_{\boldsymbol{x}}u)$ satisfies
$\begin{split}b_{\alpha}({\boldsymbol{u}},{\boldsymbol{v}})+c_{4}({\boldsymbol{u}},{\boldsymbol{v}})=\alpha\langle(f,\boldsymbol{0},u_{0})\,,\,\mathcal{G}{\boldsymbol{v}}\rangle_{L}+\langle\phi_{N}+\mathcal{V}^{-1}(\tfrac{1}{2}-\mathcal{K})u_{D}\,,\,v|_{\Sigma}\rangle_{\Sigma}\\ \text{for all }{\boldsymbol{v}}=(v,{\boldsymbol{\tau}})\in U.\end{split}$ (3.11)
By considering ${\boldsymbol{v}}\in U_{D}$ and recalling that $\mathcal{G}(U_{D})=L$, one sees that (3.11) conversely implies (3.2) and thus also (3.10) and the equivalent (3.4). Note that the left-hand side and the right-hand side in (3.11) are linear and continuous with respect to inputs ${\boldsymbol{u}},{\boldsymbol{v}}\in U$.
We want to show that the combined bilinear form $b_{\alpha}+c_{4}:U\times U\to\mathbb{R}$ is coercive for sufficiently large $\alpha>0$ and arbitrary diffusion matrix ${\boldsymbol{A}}$ in the interior. As for the other bilinear forms $b_{\alpha}+c_{i}$, $i\in\{1,2,3\}$, this directly follows from the following lemma; see Theorem 3.4.
Lemma 3.6.
The continuous bilinear form $c_{4}:U\times U\to\mathbb{R}$ is coercive on the kernel of $\mathcal{G}$, i.e.,
$c_{4}({\boldsymbol{u}},{\boldsymbol{u}})\gtrsim\|{\boldsymbol{u}}\|_{U}^{2}\quad\text{for all }{\boldsymbol{u}}\in{\rm ker}\,\mathcal{G}.$
Proof.
Let ${\boldsymbol{u}}=(u,{\boldsymbol{\sigma}})\in{\rm ker}\,\mathcal{G}$. Then, Lemma 3.1 yields the existence of a constant $C>0$ such that
$c_{4}({\boldsymbol{u}},{\boldsymbol{u}})$ $=\langle-{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}+\mathcal{V}^{-1}(\tfrac{1}{2}-\mathcal{K})(u|_{\Sigma})\,,\,u|_{\Sigma}\rangle_{\Sigma}$
$=\|{\boldsymbol{A}}^{1/2}\nabla_{\boldsymbol{x}}u\|_{L^{2}(Q)}^{2}+\tfrac{1}{2}\|u(T,\cdot)\|_{L^{2}(\Omega)}^{2}+\langle\mathcal{V}^{-1}(\tfrac{1}{2}-\mathcal{K})(u|_{\Sigma})\,,\,u|_{\Sigma}\rangle_{\Sigma}$
$\geq C\|{\boldsymbol{u}}\|_{U}^{2}+\langle\mathcal{V}^{-1}(\tfrac{1}{2}-\mathcal{K})(u|_{\Sigma})\,,\,u|_{\Sigma}\rangle_{\Sigma}.$
It remains to show that the second term is nonnegative. To this end, we consider the function
$w:=\widetilde{\mathcal{K}}(u|_{\Sigma})-\widetilde{\mathcal{V}}\phi\quad\text{for }\phi:=\mathcal{V}^{-1}(\mathcal{K}-\tfrac{1}{2})(u|_{\Sigma})$
on the exterior space-time cylinder $Q^{\rm ext}$. By restricting $w$ to $\Sigma$, we see that
$w|_{\Sigma}=(\mathcal{K}+\tfrac{1}{2})(u|_{\Sigma})-\mathcal{V}\phi=u|_{\Sigma}.$
On the other hand, from $\partial_{t}w-\Delta_{\boldsymbol{x}}w=0$, $w(0,\cdot)=0$, and the fact that $w$ satisfies the radiation condition (1.1g) (with $u^{\rm ext}$ replaced by $w$; see (2.20)), we get as in (2.18) the representation formula $w=\widetilde{\mathcal{K}}(w|_{\Sigma})-\widetilde{\mathcal{V}}(\nabla_{\boldsymbol{x}}w\cdot{\boldsymbol{n}}_{\boldsymbol{x}})$ and thus $\widetilde{\mathcal{V}}(\phi)=\widetilde{\mathcal{V}}(\nabla_{\boldsymbol{x}}w\cdot{\boldsymbol{n}}_{\boldsymbol{x}})$ and finally $\phi=\nabla_{\boldsymbol{x}}w\cdot{\boldsymbol{n}}_{\boldsymbol{x}}$. Overall, this yields that
$\langle\mathcal{V}^{-1}(\tfrac{1}{2}-\mathcal{K})(u|_{\Sigma})\,,\,u|_{\Sigma}\rangle_{\Sigma}=\langle-\phi\,,\,u|_{\Sigma}\rangle_{\Sigma}=\langle-\nabla_{\boldsymbol{x}}w\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\,,\,w|_{\Sigma}\rangle_{\Sigma}.$
Lemma 3.1 applied to ${\boldsymbol{u}}=(w,-\nabla_{\boldsymbol{x}}w)$, ${\boldsymbol{A}}={\rm Id}$, and $Q_{R}^{\rm ext}$ instead of $Q$ gives that
$\langle-\nabla_{\boldsymbol{x}}w\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\,,\,w|_{\Sigma}\rangle_{\Sigma}=\|\nabla_{\boldsymbol{x}}w\|_{L^{2}(Q_{R}^{\rm ext})}^{2}+\tfrac{1}{2}\|w(T,\cdot)\|_{L^{2}(\Omega_{R}^{\rm ext})}^{2}-\langle\nabla_{\boldsymbol{x}}w\cdot{\boldsymbol{n}}_{R,{\boldsymbol{x}}}\,,\,w|_{\Sigma_{R}}\rangle_{\Sigma_{R}}.$
Elementary analysis and the stronger radiation condition $w({\boldsymbol{x}})={\mathcal{O}}(|x|^{-d})$ and $\nabla_{\boldsymbol{x}}w({\boldsymbol{x}})\cdot{\boldsymbol{x}}={\mathcal{O}}(|x|^{-d})$ as $|x|\to\infty$ from Remark 2.3 show that $\langle\nabla_{\boldsymbol{x}}w\cdot{\boldsymbol{n}}_{R,{\boldsymbol{x}}}\,,\,w|_{\Sigma_{R}}\rangle_{\Sigma_{R}}\to 0$ as $R\to\infty$; see also the proof of [5, Theorem 3.11]. The monotone convergence theorem yields that
$\langle-\nabla_{\boldsymbol{x}}w\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\,,\,w|_{\Sigma}\rangle_{\Sigma}=\|\nabla_{\boldsymbol{x}}w\|_{L^{2}(Q^{\rm ext})}^{2}+\tfrac{1}{2}\|w(T,\cdot)\|_{L^{2}(\Omega^{\rm ext})}^{2}.$
We conclude that $\langle-\nabla_{\boldsymbol{x}}w\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\,,\,w|_{\Sigma}\rangle_{\Sigma}$ is indeed nonnegative and thus the proof. ∎
In the next theorem, we finally show that, conversely, the first component of the unique solution ${\boldsymbol{u}}=(u,{\boldsymbol{\sigma}})\in U$ of the variational formulation (3.11), which itself is equivalent to the first-order system (3.2) and the boundary integral equation (3.4), and the associated exterior function
$u^{\rm ext}:=\widetilde{\mathcal{K}}(u|_{\Sigma}-u_{D})+\widetilde{\mathcal{V}}({\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}+\phi_{N})\quad\text{on }Q^{\rm ext}$ (3.12)
are weak solutions of the transmission problem (1.1). From this and Lemma 3.6 in combination with Theorem 3.4, we immediately deduce unique solvability of (1.1).
Theorem 3.7.
If $u$ and $u^{\rm ext}$ are weak solutions of (1.1), then $(u,{\boldsymbol{\sigma}})=(u,-{\boldsymbol{A}}\nabla_{\boldsymbol{x}}u)\in U$ satisfies (3.2) and (3.4), and $u^{\rm ext}$ is given by (3.12). Conversely, if ${\boldsymbol{u}}=(u,{\boldsymbol{\sigma}})\in U$ solves (3.2) and (3.4), then $u$ and $u^{\rm ext}$ given by (3.12) are weak solutions of (1.1). Since (3.2) and (3.4) are equivalent to the (for sufficiently large $\alpha>0$) uniquely solvable variational formulation (3.11), this shows that there is indeed a unique pair $u$ and $u^{\rm ext}$ of weak solutions of (1.1).
Proof.
We have already seen that for any weak solution $u$ of (1.1), the corresponding $(u,-{\boldsymbol{A}}\nabla_{\boldsymbol{x}}u)\in U$ satisfies (3.2) and (3.4). Moreover, (3.12) is just (2.19) with $-{\boldsymbol{A}}\nabla_{\boldsymbol{x}}u$ replaced by the second variable ${\boldsymbol{\sigma}}$.
Conversely, any $(u,{\boldsymbol{\sigma}})\in U$ satisfies $u\in L^{2}(I;H^{1}(\Omega))\cap H^{1}(I;H^{-1}(\Omega))$ according to (3.1), and the first-order system (3.2) implies the heat equation in the interior (1.1a) with initial condition (1.1c). Moreover, $u^{\rm ext}$ from (3.12) satisfies the homogeneous heat equation in the exterior (1.1b) with initial condition (1.1d). By restricting $u^{\rm ext}$ to the lateral boundary $\Sigma$, we see that
$u^{\rm ext}|_{\Sigma}=({\mathcal{K}}+1/2)(u|_{\Sigma}-u_{D})+{\mathcal{V}}({\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}+\phi_{N}),$ (3.13)
and (3.4) shows the Dirichlet jump condition (1.1e), i.e., $u|_{\Sigma}-u^{\rm ext}|_{\Sigma}=u_{D}$. According to (2.20), $u^{\rm ext}$ satisfies the radiation condition (1.1g), from which we derive with the representation formula on $Q^{\rm ext}_{R}$ that $u^{\rm ext}=\widetilde{\mathcal{K}}(u^{\rm ext}|_{\Sigma})-\widetilde{\mathcal{V}}(\nabla_{\boldsymbol{x}}u^{\rm ext}\cdot{\boldsymbol{n}}_{\boldsymbol{x}})$; see Section 2.3. By restricting this representation to $\Sigma$ and exploiting $u|_{\Sigma}-u^{\rm ext}|_{\Sigma}=u_{D}$, we see that
$u^{\rm ext}|_{\Sigma}=(\mathcal{K}+1/2)(u|_{\Sigma}-u_{D})-\mathcal{V}(\nabla_{\boldsymbol{x}}u^{\rm ext}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}).$
By subtracting the latter equation from (3.13), we see that $0=\mathcal{V}(({\boldsymbol{\sigma}}+\nabla_{\boldsymbol{x}}u^{\rm ext})\cdot{\boldsymbol{n}}_{\boldsymbol{x}}+\phi_{N})$. Since $\mathcal{V}$ is an isomorphism and ${\boldsymbol{\sigma}}=-{\boldsymbol{A}}\nabla_{\boldsymbol{x}}u$, we conclude the Neumann jump condition (1.1f), i.e., $({\boldsymbol{A}}\nabla_{\boldsymbol{x}}u-\nabla_{\boldsymbol{x}}u^{\rm ext})\cdot{\boldsymbol{n}}_{\boldsymbol{x}}=\phi_{N}$, and thus the proof. ∎
Remark 3.8.
The equivalence of Theorem 3.7 follows analogously if the boundary integral equation (3.4) is replaced by the boundary integral equation (3.6), which results from the application of the exterior normal derivative to the representation (2.19) of the exterior solution $u^{\rm ext}$.
4. Numerical experiments
We present numerical experiments on the two-dimensional spatial domain $\Omega=(0,1)^{2}$ with final time $T=1$. In all experiments, ${\boldsymbol{A}}$ is the identity matrix. We use the coupling approach involving the Dirichlet trace of the representation formula from Section 3.2.1. To discretize the trial space
$U=\boldsymbol{H}(\operatorname{\mathrm{div}};Q)\cap(L^{2}(I;H^{1}(\Omega))\times L^{2}(Q)^{d}),$
we consider a partition $I_{h_{t}}$ of $I$ into intervals of size $h_{t}$ and a conforming partition $\Omega_{h_{{\boldsymbol{x}}}}$ of $\Omega$ into shape-regular triangles of size $h_{\boldsymbol{x}}$. In particular, this gives rise to a partition
$Q_{h}:=\left\{J\times K\mid J\in I_{h_{t}},K\in\Omega_{h_{{\boldsymbol{x}}}}\right\}$
of $Q=I\times\Omega$ into prisms. In time, we use the space $S^{1}(I_{h_{t}})$ of globally continuous $I_{h_{t}}$-piecewise affine functions for the first variable $u$ and the space $P^{0}(I_{h_{t}})$ of $I_{h_{t}}$-piecewise constant functions for the second variable ${\boldsymbol{\sigma}}$. In the spatial direction, we use the space $S^{1}(\Omega_{h_{{\boldsymbol{x}}}})$ of globally continuous $\Omega_{h_{{\boldsymbol{x}}}}$-piecewise affine functions for the first variable $u$ and the space $RT^{0}(\Omega_{h_{{\boldsymbol{x}}}})$ of lowest-order Raviart–Thomas functions for the second variable ${\boldsymbol{\sigma}}$. As in [27], we use the resulting tensor-product space
$U_{h}:=[S^{1}(I_{h_{t}})\otimes S^{1}(\Omega_{h_{{\boldsymbol{x}}}})]\times[P^{0}(I_{h_{t}})\otimes RT^{0}(\Omega_{h_{{\boldsymbol{x}}}})]\subset U$
as conforming approximation space of $U$. For given data $(f,u_{0},u_{D},\phi_{N})$, the discrete solution ${\boldsymbol{u}}_{h}=(u_{h},{\boldsymbol{\sigma}}_{h})\in U_{h}$ is then given by
$\begin{split}b_{\alpha}({\boldsymbol{u}}_{h},{\boldsymbol{v}}_{h})+c_{1}({\boldsymbol{u}}_{h},{\boldsymbol{v}}_{h})=\alpha\langle(f,\boldsymbol{0},u_{0})\,,\,\mathcal{G}{\boldsymbol{v}}_{h}\rangle_{L}+\langle-\mathcal{V}\phi_{N}+(\mathcal{K}-\tfrac{1}{2})u_{D}\,,\,{\boldsymbol{\tau}}_{h}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\rangle_{\Sigma}\end{split}$
for all ${\boldsymbol{v}}_{h}=(v_{h},{\boldsymbol{\tau}}_{h})\in U_{h}$. Lemma 3.2 and Theorem 3.4 yield coercivity of the bilinear form on the left-hand side provided that $\alpha$ is chosen sufficiently large, which guarantees unique solvability of this discrete variational formulation. In our experiments, we choose $\alpha=1$. The numerical eigenvalue study provided in Section 4.5 indeed suggests that this induces a coercive bilinear form.
More precisely, we consider the following concrete sequence of uniform meshes on $Q=(0,1)^{3}$: For given $\ell\in\mathbb{N}$, we first split the spatial domain $\Omega=(0,1)^{2}$ in each direction into $12^{\ell}-$ 1 uniform intervals. The resulting squares are then bisected in northwest-southeast direction to obtain a conforming shape-regular triangulation $\Omega_{h_{{\boldsymbol{x}}}}$. Finally, the time interval $I=(0,1)$ is split into $12^{\ell}-$ 1 uniform intervals, forming $I_{h_{t}}$. With the number $N_{h}=\dim(S^{1}(I_{k})\otimes S^{1}(\Omega_{h}))$ of nodes in $Q_{h}$ and the number of overall degrees of freedom $M_{h}:=\dim U_{h}$, this yields that
$h:=h_{t}\eqsim h_{{\boldsymbol{x}}}\eqsim N_{h}^{-1/3}\eqsim M_{h}^{-1/3}.$
For $\alpha$ sufficiently large such that $b_{\alpha}+c_{1}$ is coercive, the Céa lemma implies quasi-optimality of the method. That is, for the exact solution $(u,{\boldsymbol{\sigma}})\in U$ and numerical approximation $(u_{h},{\boldsymbol{\sigma}}_{h})\in U_{h}$, it holds that
$\|(u,{\boldsymbol{\sigma}})-(u_{h},{\boldsymbol{\sigma}}_{h})\|_{U}\eqsim\min_{(v_{h},{\boldsymbol{\tau}}_{h})\in U_{h}}\|(u,{\boldsymbol{\sigma}})-(v_{h},{\boldsymbol{\tau}}_{h})\|_{U}.$
For smooth exact solutions, [27] shows that
$\min_{(v_{h},{\boldsymbol{\tau}}_{h})\in U_{h}}\|(u,{\boldsymbol{\sigma}})-(v_{h},{\boldsymbol{\tau}}_{h})\|_{U}=\mathcal{O}(h).$
The experiments are implemented in MATLAB. For the calculation of the discrete boundary integral operators, we have analytically integrated integrals in time, cf. [5, 43, 28], which can be done up to evaluations of exponential integrals. The latter can be evaluated directly in MATLAB. The remaining integrals in space are calculated either by tensor Gauss–Legendre rules or adapted quadrature rules from [47] in case of (logarithmic) singularities. All systems are solved by the MATLAB backslash operator without preconditioning.
4.1. Smooth solutions vanishing in the exterior domain
We prescribe an exact interior and exterior solution as
$u(t,{\boldsymbol{x}})$ $=G(t,{\boldsymbol{x}}-{\boldsymbol{x}}^{\rm ext}),$
$u^{\rm ext}(t,{\boldsymbol{x}})$ $=0,$
with ${\boldsymbol{x}}^{\rm ext}=(-0.5,-0.5)\in\Omega^{\rm ext}$, and compute the data thereof. In particular, $f=0$, $u_{0}=0$, and $u_{D}$ and $\phi_{N}$ are smooth. On a uniform sequence of meshes with $N_{h}$ nodes, we expect
$\|(u,{\boldsymbol{\sigma}})-(u_{h},{\boldsymbol{\sigma}}_{h})\|_{U}^{2}={\mathcal{O}}(h^{2})={\mathcal{O}}(N_{h}^{-2/3}).$
In Figure 4.1, we plot the (squared) errors
$\|\operatorname{\mathrm{div}}({\boldsymbol{u}}-{\boldsymbol{u}}_{h})\|_{L^{2}(Q)}^{2},\qquad\|u-u_{h}\|_{L^{2}(Q)}^{2},\qquad\|\nabla_{\boldsymbol{x}}(u-u_{h})\|_{L^{2}(Q)}^{2},\qquad\|{\boldsymbol{\sigma}}-{\boldsymbol{\sigma}}_{h}\|_{L^{2}(Q)}^{2}.$
It is reasonable to expect higher rates for the lower-order term $\|u-u_{h}\|_{L^{2}(Q)}^{2}={\mathcal{O}}(h^{4})={\mathcal{O}}(N_{h}^{-4/3})$ and optimal order ${\mathcal{O}}(N_{h}^{-2/3})$ for the three remaining error terms. The rates are confirmed in Figure 4.1.
$10^{2}$$10^{3}$$10^{4}$$10^{-6}$$10^{-5}$$10^{-4}$$10^{-3}$2/3232/32 / $34/$ 3number of nodes $N_{h}$$\|\operatorname{\mathrm{div}}({\boldsymbol{u}}-{\boldsymbol{u}}_{h})\|_{L^{2}(Q)}^{2}$$\|u-u_{h}\|_{L^{2}(Q)}^{2}$$\|\nabla_{\boldsymbol{x}}(u-u_{h})\|_{L^{2}(Q)}^{2}$$\|{\boldsymbol{\sigma}}-{\boldsymbol{\sigma}}_{h}\|_{L^{2}(Q)}^{2}$
Figure 4.1. Smooth solution vanishing in the exterior domain.
4.2. Smooth solutions vanishing in the interior domain
We prescribe an exact interior and exterior solution as
$u(t,{\boldsymbol{x}})$ $=0,$
$u^{\rm ext}(t,{\boldsymbol{x}})$ $=G(t,{\boldsymbol{x}}-{\boldsymbol{x}}^{\rm int}),$
with ${\boldsymbol{x}}^{\rm int}=(0.5,0.5)\in\Omega$, and compute the data thereof. In particular, $f=0$, $u_{0}=0$, and $u_{D}$ and $\phi_{N}$ are smooth. In Figure 4.2, we plot the same errors as in Section 4.1, expecting similar rates.
$10^{2}$$10^{3}$$10^{4}$$10^{-6}$$10^{-5}$$10^{-4}$$10^{-3}$$10^{-2}$2/3232/32 / $34/$ 3number of nodes $N_{h}$$\|\operatorname{\mathrm{div}}({\boldsymbol{u}}-{\boldsymbol{u}}_{h})\|_{L^{2}(Q)}^{2}$$\|u-u_{h}\|_{L^{2}(Q)}^{2}$$\|\nabla_{\boldsymbol{x}}(u-u_{h})\|_{L^{2}(Q)}^{2}$$\|{\boldsymbol{\sigma}}-{\boldsymbol{\sigma}}_{h}\|_{L^{2}(Q)}^{2}$
Figure 4.2. Smooth solution vanishing in the interior domain.
4.3. Smooth solutions
We prescribe an exact interior and exterior solution as
$u(t,{\boldsymbol{x}})$ $=x_{1}^{2}+t,$
$u^{\rm ext}(t,{\boldsymbol{x}})$ $=G(t,{\boldsymbol{x}}-{\boldsymbol{x}}^{\rm int}),$
with ${\boldsymbol{x}}^{\rm int}=(0.5,0.5)\in\Omega$, and compute the data thereof. In particular, $f=-1$, $u_{0}({\boldsymbol{x}})=x_{1}^{2}$, and $u_{D}$ and $\phi_{N}$ are smooth. In Figure 4.3, we plot the same errors as in Section 4.1, expecting similar rates.
$10^{2}$$10^{3}$$10^{4}$$10^{-6}$$10^{-5}$$10^{-4}$$10^{-3}$$10^{-2}$$10^{-1}$2/3232/32 / $34/$ 3number of nodes $N_{h}$$\|\operatorname{\mathrm{div}}({\boldsymbol{u}}-{\boldsymbol{u}}_{h})\|_{L^{2}(Q)}^{2}$$\|u-u_{h}\|_{L^{2}(Q)}^{2}$$\|\nabla_{\boldsymbol{x}}(u-u_{h})\|_{L^{2}(Q)}^{2}$$\|{\boldsymbol{\sigma}}-{\boldsymbol{\sigma}}_{h}\|_{L^{2}(Q)^{2}}^{2}$
Figure 4.3. Smooth solution.
4.4. Nonsmooth solutions
We use the data $f,u_{D},\phi_{N}=0$ and $u_{0}=1$. The pair of exact solutions $(u,u^{\rm ext})$ is unknown. Since the initial conditions do not match with the Dirichlet transmission conditions, we expect singularities in the solutions leading to sub-optimal convergence rates. In Figure 4.4, we plot
$\|\mathcal{G}({\boldsymbol{u}}-{\boldsymbol{u}}_{h})\|_{L}^{2}=\|(f,0,u_{0})-\mathcal{G}{\boldsymbol{u}}_{h}\|_{L}^{2},$
which constitutes a computable lower bound of the (squared) error $\|{\boldsymbol{u}}-{\boldsymbol{u}}_{h}\|_{U}^{2}$. We observe that this error measure converges approximately with rate $N_{h}^{-1/5}$.
$10^{2}$$10^{3}$$10^{4}$$10^{-1.6}$$10^{-1.4}$$10^{-1.2}$$10^{-1}$2/3232/32 / 31/5151/51 / 5number of nodes $N_{h}$$\|\mathcal{G}({\boldsymbol{u}}-{\boldsymbol{u}}_{h})\|_{L}^{2}$
Figure 4.4. Nonsmooth solution.
4.5. Dependence of coercivity on $\boldsymbol{\alpha}$
In this section, we investigate the influence of $\alpha$ on the coercivity of the bilinear form $b_{\alpha}+c_{1}$. If $\alpha$ is sufficiently large, Lemma 3.2 and Theorem 3.4 guarantee the existence of a constant $C_{\text{\rm coer}}(\alpha)>0$ such that
$b_{\alpha}({\boldsymbol{v}},{\boldsymbol{v}})+c_{1}({\boldsymbol{v}},{\boldsymbol{v}})\geq C_{\text{\rm coer}}(\alpha)\|{\boldsymbol{v}}\|_{U}^{2}\quad\text{for all }{\boldsymbol{v}}\in U.$ (4.1)
For the discrete subspace $U_{h}$ of dimension $M_{h}$, let $\textbf{A}_{\alpha,h}\in\mathbb{R}^{M_{h}\times M_{h}}$ be the system matrix of the bilinear form $b_{\alpha}+c_{1}$, and $\textbf{U}_{h}\in\mathbb{R}^{M_{h}\times M_{h}}$ be the (symmetric and positive definite) system matrix of the inner product on $U$ with respect to some fixed basis of $U_{h}$. With the symmetric part $\textbf{A}_{\alpha,h}^{\rm s}:=(\textbf{A}_{\alpha,h}+\textbf{A}_{\alpha,h}^{\top})/2$ of $\textbf{A}_{\alpha,h}$, it holds that
$\inf_{{\boldsymbol{v}}_{h}\in U_{h}}\frac{b_{\alpha}({\boldsymbol{v}}_{h},{\boldsymbol{v}}_{h})+c_{1}({\boldsymbol{v}}_{h},{\boldsymbol{v}}_{h})}{\|{\boldsymbol{v}}_{h}\|_{U}^{2}}=\inf_{\textbf{y}\in\mathbb{R}^{M_{h}}}\frac{\textbf{y}^{\top}\textbf{A}_{\alpha,h}\textbf{y}}{\textbf{y}^{\top}\textbf{U}_{h}\textbf{y}}=\inf_{\textbf{y}\in\mathbb{R}^{M_{h}}}\frac{\textbf{y}^{\top}\textbf{A}_{\alpha,h}^{\rm s}\textbf{y}}{\textbf{y}^{\top}\textbf{U}_{h}\textbf{y}}=\inf_{\textbf{y}\in\mathbb{R}^{M_{h}}}\frac{\textbf{y}^{\top}\textbf{U}_{h}^{-1/2}\textbf{A}_{\alpha,h}^{\rm s}\textbf{U}_{h}^{-1/2}\textbf{y}}{\textbf{y}^{\top}\textbf{y}}.$
As $\textbf{U}_{h}^{-1/2}\textbf{A}_{\alpha,h}^{\rm s}\textbf{U}_{h}^{-1/2}$ is symmetric, the infimum on the right-hand side is equal to the smallest eigenvalue of this matrix. Hence, we consider the equivalent generalized eigenvalue problem
$\textbf{A}_{\alpha,h}^{\rm s}\textbf{y}=\lambda\textbf{U}_{h}\textbf{y}.$
In Table 1, we provide the smallest eigenvalue in function of $\alpha$ (columns) and meshes of different size (rows).
$\𝛼N_{h}\textbackslash\alpha$ 10000 1000 100 10 1 0.1 0.01 0.001 0.0001
64 89.1751 10.0269 1.6216 0.6109 0.0593 -0.7387 -1.1230 -1.1641 -1.1682
512 22.3626 3.0514 0.7684 0.3073 0.0223 -0.7546 -1.1696 -1.2150 -1.2195
4096 6.4107 1.1080 0.5000 0.2201 0.0124 -0.7617 -1.1906 -1.2378 -1.2426
32768 NaN 0.6609 0.2917 0.1907 0.0091 -0.7664 -1.2072 -1.2561 -1.2611
Table 1. Minimum of $\frac{b_{\alpha}({\boldsymbol{v}}_{h},{\boldsymbol{v}}_{h})+c_{1}({\boldsymbol{v}}_{h},{\boldsymbol{v}}_{h})}{\|{\boldsymbol{v}}_{h}\|_{U}^{2}}$ over all ${\boldsymbol{v}}_{h}\in U_{h}$.
We see that for $\alpha\in\left\{10000,1000,100,10,1\right\}$, all eigenvalues are positive and appear to be bounded from below by some positive constant. For $\alpha\in\left\{0.1,0.01,0.001,0.0001\right\}$, the smallest eigenvalue is already negative.
Acknowledgement
TF acknowledges funding by ANID Chile through FONDECYT project 1210391. GG acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813. MK acknowledges funding by ANID Chile through FONDECYT project 1210579.